# DiLA: Enhancing LLM Tool Learning with Differential Logic Layer

### v3

@@ Abstract
Considering the challenges faced by large language models (LLMs) in logical reasoning and planning, prior efforts have sought to augment LLMs with access to external solvers. While progress has been made on simple reasoning problems, solving classical constraint satisfaction problems, such as the Boolean Satisfiability Problem (SAT) and Graph Coloring Problem (GCP), remains difficult for off-the-shelf solvers due to their intricate expressions and exponential search spaces. In this paper, we propose a novel differential logic layer-aided language modeling (DiLA) approach, where logical constraints are integrated into the forward and backward passes of a network layer, providing another option for LLM tool learning. In DiLA, LLM aims to transform the language description to logic constraints and identify initial solutions of the highest quality, while the differential logic layer focuses on iteratively refining the LLM-prompted solution. Leveraging the logic layer as a bridge, DiLA enhances the logical reasoning ability of LLMs on a range of reasoning problems encoded by Boolean variables, guaranteeing the efficiency and correctness of the solution process. We evaluate the performance of DiLA on three classic constraint satisfaction problems and empirically demonstrate its consistent outperformance against existing prompt-based and solver-aided approaches.
DiLA: Enhancing LLM Tool Learning with Differential Logic Layer
Yu Zhang1  Hui-Ling Zhan2  Zehua Pei1  Yingzhao Lian2  Lihao Yin2 Mingxuan Yuan2Bei Yu1 1The Chinese University of Hong Kong  2Noah’s Ark Lab, Huawei
1 Introduction
@@
Recently, a significant research thrust has been on leveraging large language models (LLMs) for reasoning and planning, with numerous efforts aimed at augmenting their reasoning capabilities. These endeavors include text-based temporal reasoning (Xiong et al., 2024), logic feedback-enhanced alignment methods (Nguyen et al., 2023), and prompt-based methods such as the chain of thoughts (CoT) (Wang et al., 2023) or a simple directive like “Let’s think step by step” (Lightman et al., 2023). Despite the strides made by LLMs in achieving human-like reasoning abilities, they still encounter challenges when confronted with complex plannings (Valmeekam et al., 2022). Occasionally, LLMs exhibit unfaithful reasoning, leading to derived conclusions that do not consistently follow the previously generated reasoning chain in practical applications (Pan et al., 2023).
@@
To this end, recent works have begun to augment LLMs with access to external solvers, i.e., utilize LLMs to first parse natural language logical questions into symbolic representations and subsequently employ external solvers to generate answers based on these representations. To enhance parsing accuracy, LoGiPT (Feng et al., 2023) has been proposed to directly emulate the reasoning processes and mitigate the parsing errors by learning to strict adherence to solver syntax and grammar. It is fine-tuned on a constructed instruction-tuning dataset derived from revealing and refining the invisible reasoning process of deductive solvers. An alternative approach for improving reasoning capabilities involves LOGIC-LM, a new symbolic solver-aided modeling approach (Pan et al., 2023), in which an LLM is used to generate a symbolic problem formulation with in-context learning and offload the actual reasoning task to off-the-shelf symbolic reasoners, such as Z3 solver (De Moura and Bjørner, 2008). This solver-augmented method has demonstrated superiority over CoT by achieving a large margin of $18.4\%$ on average, establishing a new state-of-the-art.
@@
Despite their impressive performance on various benchmark tests, LOGIC-LM (Pan et al., 2023) and its solver-augmented successors can not deal with reasoning and planning problems in practical scenarios. Particularly, these real problems often state a set of premises and complex constraints and require a sophisticated search process to find the optimal solution, which is still challenging even for modern solvers. More specifically, even the state-of-the-art symbolic solvers, such as Z3(De Moura and Bjørner, 2008) and Kissat(Biere and Fleury, 2022), encounter significant bottlenecks when addressing formal verification problems such as electronic circuit verification involving tree/cyclic circuit structures(Shi et al., 2023) and job-shop scheduling problems (Li et al., 2022) with complex graph connections. For instance, the Kissat solver requires days or even weeks to process a vanilla circuit verification problem with arithmetic circuit modules including multipliers or multiply-add circuits, yet fails to produce a solution. This highlights the limitations of solver-augmented LLM approaches which solely rely on symbolic solvers to tackle logical reasoning problems in reality.
In this paper, we propose DiLA, a novel tool-learning approach for Large Language Models, designed to enhance their logical reasoning capabilities through the integration of an additional logic layer. Unlike existing methods that either rely on in-context prompting for step-by-step reasoning (see Figure 1(left)) or entirely offload reasoning to external solvers (see Figure 1(middle)), our approach offers a third option: enhancing LLMs’ reasoning ability by incorporating a differential logic layer (see Figure 1(right)). Specifically, DiLA leverages the LLM to parse and comprehend the problem description, generate an initial solution based on its language understanding, and then iteratively refine this solution through forward and backward passes of a network layer—the differential logic layer—which embeds first-order logic constraints into its architecture. In this way, DiLA overcomes the limitations of traditional solvers by directly performing reasoning within the framework of layer-augmented LLMs. Our contributions are summarized as follows:
@@
  • Compared to other solver-aided LLM approaches, DiLA pioneers a novel strategy that differentiates symbolic problems and iteratively searches for solutions through forward and backward propagation of a network layer, thereby circumventing the limitations of off-the-shelf solvers.
@@
We evaluate the performance of our approach on three constraint satisfaction problems: logical deduction, Boolean satisfiability, and graph coloring. Our analysis yields two key findings: firstly, on simple artificial problem instances, DiLA boosts the inference accuracy of LLMs to 100% and consistently outperforms solver-aided approaches with improved runtime. Secondly, for real-world problems that current solvers struggle with, DiLA showcases robustness and remarkable efficiency in handling these complex test cases, thereby opening up opportunities for further real-world applications.
2 Motivation
@@
Figure 2: Exemplar comparison of solving graph coloring problems by different approaches. Direct prompts by GPT-4 make errors when generating the color assignment step by step; LOGIC-LM based on the Z3 symbolic solver cannot solve parsed constraints and outputs unknown (both errors are highlighted in red). In contrast, DiLA can generate the correct answer by combining the strengths of LLMs and the differential logic layer.
@@
This paper explores the potential of language understanding and logical reasoning capabilities in LLMs. Traditionally, the prevailing approaches have either relied solely on LLMs for step-by-step reasoning (Wang et al., 2023; Zhou et al., 2022) or offloaded reasoning tasks to off-the-shelf solvers (Olausson et al., 2023; Ye X, 2023). However, we propose a third approach, as these two extremes either underutilize or over-rely on LLMs during reasoning. Specifically, DiLA leverages the powerful understanding abilities of LLMs to extract logical formulas and generate possible solutions. By utilizing LLMs, such as GPT-4, as a solution generation engine, we can produce an initial solution based on input semantic constraints, laying the groundwork for further refinement.
@@
The second motivation for this research stems from an analysis of current solver applications in real-world scenarios, characterized by two distinctive features: (1) The large scale of reasoning problems in reality, leading to rapid degradation in the performance of heuristic-based solvers due to the exponential expansion of the search space; (2) The formidable challenge presented by the complex structure of logic formulas, often requiring weeks or even months to resolve using current solvers. Both characteristics significantly limit the effectiveness of a solver-augmented LLM in the domain of logic reasoning.
@@
Consequently, our objective is to identify a synergistic approach that combines the strengths of LLMs and a differential logic layer, thereby circumventing the limitations of traditional solvers. This approach leverages the LLM’s capacity to comprehend logical formulas while concurrently utilizing the logic layer’s refinement abilities to achieve accurate solutions. Figure 2 illustrates a comparison between CoT, LOGIC-LM, and our proposed DiLA. The LLM alone may introduce logical flaws during step-by-step inference, such as assigning the same color to vertex 1 and vertex 4 despite their edge connection, and LOGIC-LM may struggle with complex reasoning problems due to its backbone solver’s limitations. In contrast, DiLA produces accurate answers through the collaboration of the LLM and the differential logic layer. Specifically, DiLA first uses the LLM to parse a natural language input into logic constraints and generate an initial solution based on its semantic understanding, then employs the differential logic layer to refine this initial solution.
3 Differenial Logic Layer-Aided Language Models
3.1 Overview
Figure 3: The overall flow of DiLA. Starting from a problem in natural language, in Step 1, the LLM parses it into a SAT problem specification which consists of a set of logical formulas. Then in Step 2, the backbone LLM tries to output an initial solution based on its language understanding. In Step 3, the relaxed variables and extracted constraints are offloaded to a differential logic layer, checking the constraint satisfiability, and updating the solution until until all constraints are met. Here, the solid arrow indicates the forward pass, while the dashed arrow represents the backward pass.
@@ 3.1 Overview
In this section, we present DiLA, which augments LLM with the ability of logical reasoning by incorporating a differential logic layer. More specifically, DiLA addresses the challenge of using LLMs to tackle canonical reasoning tasks expressed in natural language. These tasks typically involve presenting a set of premises and constraints, prompting questions that necessitate intricate deductive reasoning over the provided inputs, which remains a formidable challenge even for contemporary LLMs (Valmeekam et al., 2022).
@@
The general procedure for solving natural language reasoning tasks with DiLA can be conceptualized in three distinct steps: parsing, initialization, and refinement. Given a natural language input that describes both the propositional constraints $\phi$ and the question $Q$, we first parse this input into a SAT specification using LLMs, thereby obtaining a formal description of the constraints and variables. Next, we leverage the LLM’s natural language understanding to generate an initial variable assignment. Since this initial solution may only partially satisfy the propositional constraints, DiLA iteratively refines it through a differential logic layer that encodes all of the logical formulas, ultimately yielding a more accurate solution. Figure 3 illustrates the overall flow of our proposed DiLA.
3.2 Problem Formulator and Initialization
@@
Intuitively, LLMs may struggle with directly solving complex reasoning problems. However, they have demonstrated a notable ability to comprehend textual inputs and translate them into formal programs, such as mathematical equations (He-Yueya et al., 2023) or satisfiability modulos (Ye X, 2023). Notably, the SAT problem can serve as a versatile intermediate step for solving a broad range of constraint satisfaction problems, provided they can be expressed using Boolean variables. Specifically, problem instances from NP-complete domains, such as Graph Coloring and Set Cover, can be seamlessly encoded into SAT problem specifications, thereby allowing for efficient solutions via SAT algorithms (Stechly et al., 2024; Liu et al., 2023). Therefore, we harness SAT encoding as a general bridge to tackle these constraint satisfaction problems in practical settings.
@@
Specifically, given a problem description in natural language, DiLA prompts an LLM with detailed instructions to generate the SAT specification, which includes a set of premises and constraints. Typically, the SAT specification here involves conjunctive normal forms (CNFs), denoted as $\phi(v_{1},\dots,v_{n})$, which is a conjunction of clauses (constraints) $C$. Formally, rules are written in the conjunctive form of clauses $C_{1}\wedge C_{2}\cdots\wedge C_{m}$, which each $C_{j}$ is a constraint. The total rule is satisfied if and only if all of the clauses $C_{1},C_{2},\dots,C_{m}$ are simultaneously True. Each clause represents a disjunction of literals, where a literal is either a propositional variable $v_{i}$ or its complement $\neg v_{i}$, for example, $v_{1}\vee\neg v_{2}$. In other words, if at least one literal in a clause is True, this clause would also be True. Variables can be assigned logic values, either $1$ or $-1$, representing True or False, respectively  [^1] In other works, they may claim the logic value of each literal is 00 or $1$. It should be noted that the two claims are equal under simple mathematical transformations.. See Appendix A for more details on the SAT specification.
@@
Aside from problem formulation, leveraging an LLM can be a valuable strategy to generate an initial solution. We observe that, after generating the SAT specification, we can prompt LLMs, like GPT-4 (Achiam et al., 2023) or Llama-3 (Meta, 2024), to produce a potential solution or a set of possible solutions. This can be achieved by framing the problem as a natural language query, such as “What is the logical solution based on the premises?” or “Can you provide a possible answer from these given constraints?” (see Appendix B for more details). The LLM’s response can then serve as a starting point for further refinement and validation, allowing us to build upon its output and iteratively improve the solution through logical analysis and reasoning. By harnessing the LLM’s ability to understand semantic constraints, we can tap into its potential to facilitate the initial solution-finding process and accelerate solving progress towards a well-reasoned answer.
3.3 From SAT to Differential MaxSAT Formulations
@@
In fact, SAT is very efficient at expressing constraint satisfaction problems in that many of the standard NP-complete problems, like graph coloring, can be straightforwardly formulated and efficiently solved as SAT problems (Buss and Nordström, 2021; Liu et al., 2023). Traditionally, a SAT solver, such as Z3 and Kissat, is leveraged to determine a satisfying assignment for the given constraint formula $\phi$. Contemporary SAT solvers are founded on the Conflict-Driven-Clause-Learning (CDCL) algorithm, which excels in its ability to learn from conflicts and use that conflict knowledge to prune the branch-and-bound search space more effectively. However, existing CDCL-based SAT solvers still suffer from exponential searching space and are unable to correct errors through a learning-from- mistakes system, resulting in an infinite loop in solving complex SAT problems (Shi et al., 2023).
@@
In this study, when an NL reasoning problem is fed into DiLA, our goal is to determine its solution that can satisfy all logical constraints. To achieve this, we might incorporate a solver as an additional logic reasoner for the LLM, as in SATLM (Ye X, 2023) or LOGIC-LM (Pan et al., 2023). However, the exponential search complexity inherent in heuristic-based SAT solvers poses a significant challenge, limiting their effectiveness when dealing with complex real-world problems. Therefore, a key issue is how to design an efficient solver surrogate that can both be seamlessly integrated into LLMs and efficiently address logical reasoning problems.
@@
The maximum satisfiability (MaxSAT) problem serves as the optimization counterpart to the SAT problem, aiming to maximize the number of satisfied clauses. Indeed, if a solution to the MaxSAT problem can satisfy all the clauses, the variable assignment can be used to constitute a valid solution for the original SAT problem. In the case of the SAT problem, each CNF is associated with a set of clauses (constraints), and each clause is defined on a subset of variables, signifying the variables’ simultaneous legal assignments. Formally, each CNF $\phi(v_{1},\dots,v_{n})$ comprises $n$ binary variables, with each $v_{i}\in\{1,-1\}$ ($i\in{1,2,\dots,n}$) representing a boolean variable. Let’s introduce the coefficient $\boldsymbol{c}_{j}\in\{-1,0,1\}^{m}$, where $c_{ij}$ denotes the sign of $v_{i}$ in clause $j\in{1,2,\dots,m}$. Consequently, we can establish a clause matrix $\boldsymbol{C}\in\{1,-1,0\}^{m\times n}$, where each element $c_{ij}$ in $\boldsymbol{C}$ signifies the sign of variable $\tilde{v}_{i}$ in clause $j$. Therefore, each SAT instance can be translated into a corresponding MaxSAT problem, wherein $\bigvee$ represents the logical “or” symbol,
@@
where $\bigwedge$ is the logical “and” symbol. Indeed, the objective value in Equation 2 is 0 if and only if a satisfiable solution can be found. Our goal is to establish a continuous upper bound, referred to as the “loss”, for each clause to quantify its level of unsatisfiability. In essence, the loss takes an upper bound if the clause is unsatisfied, and by minimizing this loss, we can strive to push it closer to satisfaction. To make a purely quadratic loss function as in  (Wang and Kolter, 2019), we introduce $v_{0}=1$ and $s_{0j}=-1$ in Equation 2. Therefore, the minimization problem in Equation 2 can be solved by transforming into a quadratic loss function as:
@@
Figure 4: Examplar illustration of VCG.
3.4 Differential Logic Layer
We envision the logic layer being used primarily at the top of LLMs, embedding logical formulas $\phi$ produced by the backbone LLM, taking LLM-initialized variable assignments as inputs, and producing outputs that are consistent with $\phi$. Specifically, we draw an analogy between the fully connected (FC) layer and the variable-clause graph (VCG), as depicted in Figure 4, where the weight for positive literals in the clause is 1 (solid arrow) and negative literals in the clause is -1 (dashed arrow). Specifically, we map each variable to an input neuron in an FC layer, each clause to an output neuron, and the coefficients to the weights in the linear transformation with a zero bias vector. While this analogy holds, there are two crucial differences between traditional FC layers and our proposed logic layer. Firstly, the differential logic layer has no unknown parameters, whereas FC layers require data-driven training to learn their weights. Secondly, each clause in the logic layer is only partially connected to variables, in contrast to fully connected layers, where all input neurons are connected to all output neurons.
@@ 3.4 Differential Logic Layer
0:  $y^{k}$, final solution $\boldsymbol{v}^{*}$.
1:  $\tilde{\boldsymbol{v}}^{k}\leftarrow[\boldsymbol{v}^{k}_{i}>0:i=1,\dots,n]$;
2:  $\phi^{\prime}\leftarrow\phi(\tilde{v}^{k}_{1},\dots,\tilde{v}^{k}_{n})$;

### v4

@@ Abstract
[…]
Introduction
@@
Recently, a significant research thrust has been on leveraging large language models (LLMs) for reasoning and planning, with numerous efforts aimed at augmenting their reasoning capabilities. These endeavors include text-based temporal reasoning (Xiong et al. 2024), logic feedback-enhanced alignment methods (Nguyen, Fungwacharakorn, and Satoh 2023), and prompt-based methods such as the chain of thoughts (CoT) (Wang et al. 2023a) or a simple directive like “Let’s think step by step” (Lightman et al. 2023). To further bolster the reasoning ability of LLMs, an AlphaZero-like tree-search learning framework has been introduced. This framework demonstrates how tree-search, coupled with a learned value function, can guide LLMs’ decoding and enhance their reasoning abilities (Feng et al. 2023b). Despite the strides made by LLMs in achieving human-like reasoning abilities, they still encounter challenges when confronted with complex plannings (Valmeekam et al. 2022). Occasionally, LLMs exhibit unfaithful reasoning, leading to derived conclusions that do not consistently follow the previously generated reasoning chain in practical applications (Pan et al. 2023).
@@
To this end, a growing trend in this area involves augmenting LLMs with access to external engines, i.e., utilizing LLMs to first parse natural language logical questions into symbolic representations and subsequently employ external solvers to generate answers based on these representations. To enhance parsing accuracy, LoGiPT (Feng et al. 2023a) has been proposed to directly emulate the reasoning processes and mitigate the parsing errors by learning to strict adherence to solver syntax and grammar. It is fine-tuned on a constructed instruction-tuning dataset derived from revealing and refining the invisible reasoning process of deductive solvers. An alternative approach for improving reasoning capabilities involves SATLM, a new satisfiability-aided modeling approach (Ye X 2023), in which an LLM is used to generate a declarative task specification with few-shot prompting and offload the actual reasoning task to an off-the-shelf theorem prover, i.e., Z3 solver (De Moura and Bjørner 2008). This method has demonstrated superiority over program-aided Language Models (Gao et al. 2023) by achieving a $23\%$ improvement on the GSM arithmetic reasoning dataset, establishing a new state-of-the-art.
@@
Despite their impressive performance on various benchmark tests, SATLM (Ye X 2023) and its solver-augmented successors can not deal with reasoning and planning problems in practical scenarios. Particularly, these real problems often state a set of premises and complex constraints and require a sophisticated search process to find the optimal solution, which is still challenging even for modern solvers. More specifically, even the state-of-the-art symbolic solvers, such as Z3(De Moura and Bjørner 2008) and Kissat(Biere and Fleury 2022), encounter significant bottlenecks when addressing formal verification problems such as electronic circuit verification involving tree/cyclic circuit structures(Shi et al. 2023) and job-shop scheduling problems (Li et al. 2022) with complex graph connections. For instance, the Kissat solver requires days or even weeks to process a vanilla circuit verification problem with arithmetic circuit modules including multipliers or multiply-add circuits, yet fails to produce a solution. This highlights the limitations of solver-augmented LLM approaches which solely rely on symbolic solvers to tackle logical reasoning problems in reality.
In this paper, we propose DiLA, a novel tool-learning approach for Large Language Models, designed to enhance their logical reasoning capabilities through the integration of an additional logic layer. Unlike existing methods that either rely on in-context prompting for step-by-step reasoning (see Figure 1(left)) or entirely offload reasoning to external solvers (see Figure 1(middle)), our approach offers a third option: enhancing LLMs’ reasoning ability by incorporating a differential logic layer (see Figure 1(right)). Specifically, DiLA leverages the LLM to parse and comprehend the problem description, generate an initial solution based on its language understanding, and then iteratively refine this solution through forward and backward passes of a network layer that embeds first-order logic constraints into its architecture. In this way, DiLA overcomes the limitations of traditional solvers by directly performing reasoning within the framework of layer-augmented LLMs. Our contributions are summarized as follows:
@@
  • We propose a novel neural network layer, termed the logic layer, which differentiates symbolic problems and iteratively searches for solutions through forward and backward propagation of a network layer, thereby circumventing the limitations of off-the-shelf solvers.
@@
We evaluate the performance on three constraint satisfaction problems: logical deduction, Boolean satisfiability, and graph coloring. Our analysis yields two key findings: firstly, on simple problem instances, DiLA boosts the inference accuracy of LLMs to 100% and consistently outperforms solver-aided approaches with improved runtime. Secondly, for real-world problems that current solvers struggle with, DiLA showcases robustness and remarkable efficiency in handling these complex test cases, thereby opening up opportunities for further real-world applications.
Preliminary and Related Works
Preliminary
In general, Boolean formulae are represented in Conjunctive Normal Form (CNF) as a conjunction of clauses, where a clause is a disjunction of literals and a literal denotes either a variable or its negation. Each variable can be assigned a logic value, either 0 or $1$. Any general Boolean problem can be represented as a CNF formula model. An SAT solver either finds an assignment such that CNF is satisfied or proves that no such assignment exists, i.e., UNSAT. Modern SAT solvers are based on the conflict-driven-clause-learning (CDCL) algorithm, and they work as a basic engine for many applications. SAT plays an important role in data mining applications, e.g., Maximal Frequent Subgraph Mining (Liu and Li 2022), Graph Coloring (Velev 2007), Sequence Mining (Jabbour, Sais, and Salhi 2013).
@@
Figure 2: Exemplar comparison of solving graph coloring problems by different approaches. Direct prompts by GPT-4 make errors when generating the color assignment step by step; SATLM, based on the Z3 symbolic solver, cannot solve parsed constraints and outputs unknown (both errors are highlighted in red). In contrast, DiLA can generate the correct answer by combining the strengths of LLMs and the differential logic layer.
Related Works
Logic Reasoning in LLMs. Prior approaches to NL-based reasoning with LLMs can be broadly categorized into two groups. One is in-context learning approaches that design special prompts to elicit LLMs’ step-by-step reasoning capabilities. Typical methods include chain-of-thought prompting (Wang et al. 2023a) that generates a sequence of reasoning steps before the final answer and the least-to-most prompting (Zhou et al. 2022) that breaks the problem down into simpler components that can be solved sequentially. Both the above approaches perform reasoning directly over natural language (NL), providing greater flexibility than symbolic-based reasoning. However, the intrinsic complexity and ambiguity of NL also bring undesired issues such as unfaithful reasoning and hallucinations. The other is tool-augmented approaches that only require LLMs to parse the problem specification out of the language description accurately and leverage off-the-shelf automated tools to derive the final answer, as exemplified by SATLM (Ye X 2023) and LOGIC-LM (Pan et al. 2023). The tool-augmented approaches guarantee the correctness of the answer with respect to the parsed specification and avoid planning errors in the solving process. However, the performance of such tool-augmented methods highly relies on external tools and can not deal with many real problems due to the deficiency of backbone solvers.
Neural Symbolic Reasoning. Deep neural networks have made remarkable progress in various domains, but their ability to handle logical reasoning tasks remains uncertain. One potential approach is to encode logical constraints that are essential for certain inference tasks symbolically, making them available to the model either during training or inference. SATNet (Wang et al. 2019), for instance, introduces a differentiable MaxSAT solver layer into its network architecture to learn the logical structure of problems through data-driven learning. Their approximation is based on a coordinate descent approach for solving the semidefinite program (SDP) relaxation of the MaxSAT problem. SATNet does not assume that the logical structure of the problem is given, and instead attempts to learn it. While SATNet achieves high accuracy in solving Sudoku problems, it cannot be generalized to regular SAT problems as it relies on a specific number of accurate assignments as inputs and is only applicable to solving logic problems with constant constraints after training. In the pursuit of neuro-symbolic learning, SMTLayer (Wang et al. 2023b) incorporates a satisfiability modulo theories (SMT) solver into a DNN layer, thereby embedding problem-specific theories into DNN architectures through data-driven training. Compared to SATNet, SMTLayer demonstrates superior accuracy in logical reasoning tasks and a reduced need for training data. However, SMTLayer still requires a certain amount of training data and relies on off-the-shelf SMT solvers for reasoning, making it unsuitable for industrial problem-solving. Similar to the aforementioned neuro-symbolic approach, DiLA also aims to encode logical constraints in the network layer, making them available to the model during forward and backward passes.
Motivation
@@
This paper explores the potential of language understanding and logical reasoning capabilities in LLMs. Traditionally, the prevailing approaches have either relied solely on LLMs for step-by-step reasoning (Wang et al. 2023a; Zhou et al. 2022) or offloaded reasoning tasks to off-the-shelf solvers (Olausson et al. 2023; Ye X 2023). However, we propose a third approach, as these two extremes either underutilize or over-rely on LLMs during reasoning. Specifically, DiLA leverages the powerful understanding abilities of LLMs to extract logical formulas and generate possible solutions. By utilizing LLMs, such as GPT-4, as a solution generation engine, we can produce an initial solution based on input semantic constraints, laying the groundwork for further refinement.
@@
The second motivation stems from an analysis of current solver applications in real-world scenarios, characterized by two distinctive features: (1) The large scale of reasoning problems in reality, leading to rapid degradation in the performance of heuristic-based solvers due to the exponential expansion of the search space; (2) The formidable challenge presented by the complex structure of logic formulas, often requiring weeks or even months to resolve using current solvers. Both characteristics significantly limit the effectiveness of a solver-augmented LLM in logic reasoning.
@@
Consequently, our objective is to identify a synergistic approach that combines the strengths of LLMs and a differential logic layer, thereby circumventing the limitations of traditional solvers. This approach leverages the LLM’s capacity to comprehend logical formulas while concurrently utilizing the logic layer’s refinement abilities to achieve accurate solutions. Figure 2 illustrates a comparison between CoT, SATLM, and our proposed DiLA. The LLM alone may introduce logical flaws during step-by-step inference, such as assigning the same color to vertex 1 and vertex 4 despite their edge connection, and SATLM may struggle with complex reasoning problems due to its backbone solver’s limitations. In contrast, DiLA produces accurate answers through the collaboration of the LLM and the differential logic layer. Specifically, DiLA first uses the LLM to parse a natural language input into logic constraints and generate an initial solution based on its semantic understanding, then employs the differential logic layer to refine this initial solution.
Figure 3: The overall flow of DiLA.
Differential Logic Layer-Aided LLMs
Overview
@@
In this section, we present DiLA, which augments LLM with the ability of logical reasoning by incorporating a differential logic layer. More specifically, DiLA addresses the challenge of using LLMs to tackle canonical reasoning tasks expressed in natural language. These tasks typically involve presenting a set of premises and constraints, prompting questions that necessitate intricate deductive reasoning over the provided inputs, which remains a formidable challenge even for contemporary LLMs (Valmeekam et al. 2022).
@@
The general procedure for solving natural language reasoning tasks with DiLA can be conceptualized in three distinct steps: parsing, initialization, and refinement (as illustrated in Figure 3). Given a natural language input that describes both the propositional constraints $\phi$ and the question $Q$, we first parse this input into a SAT specification using LLMs (step 1), thereby obtaining a formal description of the constraints and variables. Next, we leverage the LLM’s natural language understanding to generate an initial variable assignment (step 2). Since this initial solution may only partially satisfy the propositional constraints, DiLA iteratively refines it through a differential logic layer that encodes all of the logical formulas (step 3), ultimately yielding a more accurate solution.
Problem Formulator and Initialization
@@
Intuitively, LLMs may struggle with directly solving complex reasoning problems. However, they have demonstrated a notable ability to comprehend textual inputs and translate them into formal programs, such as mathematical equations (He-Yueya et al. 2023) or satisfiability modulo (Ye X 2023). Notably, the SAT problem can serve as a versatile intermediate step for solving a broad range of constraint satisfaction problems, provided they can be expressed using Boolean variables. Specifically, problem instances from NP-complete domains, such as Graph Coloring and Set Cover, can be seamlessly encoded into SAT problem specifications, thereby allowing for efficient solutions via SAT algorithms (Stechly, Valmeekam, and Kambhampati 2024; Liu et al. 2023). Therefore, we harness SAT encoding as a general bridge to tackle these constraint satisfaction problems in practical settings.
@@
Specifically, given a problem description in natural language, DiLA prompts an LLM with detailed instructions to generate the SAT specification, which includes a set of premises and constraints. Typically, the SAT specification here involves conjunctive normal forms (CNFs), denoted as $\phi(v_{1},\dots,v_{n})$, which is a conjunction of clauses (constraints) $C$. Formally, rules are written in the conjunctive form of clauses $C_{1}\wedge C_{2}\cdots\wedge C_{m}$, which each $C_{j}$ is a constraint. The total rule is satisfied if and only if all of the clauses $C_{1},C_{2},\dots,C_{m}$ are simultaneously True. Each clause represents a disjunction of literals, where a literal is either a propositional variable $v_{i}$ or its complement $\neg v_{i}$, for example, $v_{1}\vee\neg v_{2}$. In other words, if at least one literal in a clause is True, this clause would also be True. Variables can be assigned logic values, either $1$ or 1-1, representing True or False, respectively [^1] In other works, they may claim the logic value of each literal is 0 or $1$. It should be noted that the two claims are equal under simple mathematical transformations..
@@
Aside from problem formulation, leveraging an LLM can be a valuable strategy to generate an initial solution. We observe that, after generating the SAT specification, we can prompt LLMs, like GPT-4 (Achiam et al. 2023) or Llama-3 (Meta 2024), to produce a potential solution or a set of possible solutions. This can be achieved by framing the problem as a natural language query, such as “What is the logical solution based on the premises?” or “Can you provide a possible answer from these given constraints?”. The LLM’s response can then serve as a starting point for further refinement and validation, allowing us to build upon its output and iteratively improve the solution through logical analysis and reasoning. By harnessing the LLM’s ability to understand semantic constraints, we can tap into its potential to facilitate the initial solution-finding process and accelerate solving progress towards a well-reasoned answer.
From SAT to Differential MaxSAT
@@
Traditionally, a SAT solver, such as Z3 and Kissat, is leveraged to determine a satisfying assignment for the given constraint formula $\phi$. Contemporary SAT solvers are founded on the CDCL algorithm, which excels in its ability to learn from conflicts and use that conflict knowledge to prune the branch-and-bound search space more effectively. However, existing CDCL-based SAT solvers still suffer from exponential searching space and are unable to correct errors through a learning-from-mistakes system, resulting in an infinite loop in solving complex SAT problems (Shi et al. 2023).
@@
In this study, when an NL reasoning problem is fed into DiLA, our goal is to determine its solution that can satisfy all logical constraints. To achieve this, we might incorporate a solver as an additional logic reasoner for the LLM, as in SATLM (Ye X 2023) or LOGIC-LM (Pan et al. 2023). However, the exponential search complexity inherent in heuristic-based SAT solvers poses a significant challenge, limiting their effectiveness when dealing with complex real-world problems. Therefore, a key issue is how to design an efficient solver surrogate that can both be seamlessly integrated into LLMs and efficiently address logical reasoning problems.
@@
The maximum satisfiability (MaxSAT) problem serves as the optimization counterpart to the SAT problem, aiming to maximize the number of satisfied clauses. Indeed, if a solution to the MaxSAT problem can satisfy all the clauses, the variable assignment can be used to constitute a valid solution for the original SAT problem. In the case of the SAT problem, each CNF is associated with a set of clauses (constraints), and each clause is defined on a subset of variables, signifying the variables’ simultaneous legal assignments. Formally, each CNF $\phi(v_{1},\dots,v_{n})$ comprises $n$ binary variables, with each $v_{i}\in\{1,-1\}$ ($i\in{1,2,\dots,n}$) representing a boolean variable. The coefficients are represented as $\boldsymbol{c}_{j}\in\{-1,0,1\}^{m}$, where $c_{ij}$ denotes the sign of $v_{i}$ in clause $j\in{1,2,\dots,m}$. Consequently, we can establish a clause matrix $\boldsymbol{C}\in\{1,-1,0\}^{m\times n}$, where each element $c_{ij}$ in $\boldsymbol{C}$ signifies the sign of variable $\tilde{v}_{i}$ in clause $j$. Therefore, each SAT instance can be translated into a corresponding MaxSAT problem, wherein $\bigvee$ represents the logical “or” symbol,
@@
where $\bigwedge$ is the logical “and” symbol. Indeed, the objective value in Equation 2 is 0 if and only if a satisfiable solution can be found. Our goal is to establish a continuous upper bound, referred to as the “loss”, for each clause to quantify its level of unsatisfiability. In essence, the loss takes an upper bound if the clause is unsatisfied, and by minimizing this loss, we can strive to push it closer to satisfaction. To make a purely quadratic loss function as in  (Wang and Kolter 2019), we introduce $v_{0}=1$ and $s_{0j}=-1$ in Equation 2. Therefore, the minimization problem in Equation 2 can be solved by transforming into a quadratic loss function as:
@@
Take a simple SAT problem with clauses $(v_{1}\vee v_{2})\wedge v_{1}$ as an illustrating example. Based on Equation 3, the loss function for this SAT problem is
$\text{loss}=\text{loss}_{1}+\text{loss}_{2}=\frac{(v_{1}+v_{2}-v_{0})^{2}-1}{8}+\frac{(v_{1}-v_{0})^{2}}{4}.$
@@
Differential Logic Layer
We envision the logic layer being used primarily at the top of LLMs, embedding logical formulas $\phi$ produced by the backbone LLM, taking LLM-initialized variable assignments as inputs, and producing outputs that are consistent with $\phi$. Specifically, we draw an analogy between the fully connected (FC) layer and the variable-clause graph (VCG), where the weight for positive literals in the clause is 1 (solid arrow) and negative literals in the clause is -1 (dashed arrow). Specifically, we map each variable to an input neuron in an FC layer, each clause to an output neuron, and the coefficients to the weights in the linear transformation with a zero bias vector. While this analogy holds, there are two crucial differences between traditional FC layers and our proposed logic layer. Firstly, the logic layer has no unknown parameters, whereas FC layers require data-driven training to learn their weights. Secondly, each clause in the logic layer is only partially connected to variables, in contrast to fully connected layers, where all input neurons are connected to all output neurons.
@@
[…]